# Fantastic Copyrighted Beasts and How (Not) to Generate Them

### v1

@@ Abstract
Recent studies show that image and video generation models can be prompted to reproduce copyrighted content from their training data, raising serious legal concerns around copyright infringement. Copyrighted characters, in particular, pose a difficult challenge for image generation services, with at least one lawsuit already awarding damages based on the generation of these characters. Yet, little research has empirically examined this issue. We conduct a systematic evaluation to fill this gap. First, we build CopyCat, an evaluation suite consisting of diverse copyrighted characters and a novel evaluation pipeline. Our evaluation considers both the detection of similarity to copyrighted characters and generated image’s consistency with user input. Our evaluation systematically shows that both image and video generation models can still generate characters even if characters’ names are not explicitly mentioned in the prompt, sometimes with only two generic keywords (e.g., prompting with “videogame, plumber” consistently generates Nintendo’s Mario character). We then introduce techniques to semi-automatically identify such keywords or descriptions that trigger character generation. Using our evaluation suite, we study runtime mitigation strategies, including both existing methods and new strategies we propose. Our findings reveal that commonly employed strategies, such as prompt rewriting in the DALL·E system, are not sufficient as standalone guardrails. These strategies must be coupled with other approaches, like negative prompting, to effectively reduce the unintended generation of copyrighted characters. Our work provides empirical grounding to the discussion of copyright mitigation strategies and offers actionable insights for model deployers actively implementing them.
@@ 1 Introduction
State-of-the-art image and video generation models demonstrate a remarkable ability for generating high-quality visual content based on free-form user inputs (Rombach et al., 2022; Betker et al., 2023; Chen et al., 2024; Li et al., 2024a; Blattmann et al., 2023; Esser et al., 2024). However, recent research has shown that generative models, including image and video generation models, are susceptible to memorizing and generating entire datapoints or concepts from their training data (Somepalli et al., 2023; Carlini et al., 2023a, b). Since some training data originates from copyrighted materials (Carlini et al., 2023b; Kumari et al., 2023), regurgitation of such content may lead to legal intellectual property liability for users and model deployers who further make use of the generated content. In particular, this liability may stem not only from verbatim generation of training data points, but generation of some copyrightable repeating motifs highly similar to those from the training data. A notable example of such copyrightable data is copyrighted characters. Several legal scholars have highlighted that these characters will pose a particularly difficult, and under-explored, challenge for image or video generation services (Sag, 2023; Henderson et al., 2023; Lee et al., 2024). At least one lawsuit in China has already resulted in liability for an image generation service that generated the copyrighted character, Ultraman (Shimbun, 2024).
In this work, we provide a study focused on a subset of copyrighted content: copyrighted characters, such as popular Intellectual Property (IP) from Disney, Nintendo, and Dreamworks.111 We study the phenomenon of copyrighted characters generation from an image-based reasoning and natural language processing perspective. Artist compensation and other downstream policy issues are also important, but the formal discussion on these topics are beyond the scope of this work. For additional legal background on the distinction between copyrighted characters and other types of memorization studies, see Appendix A. Given the legal risks involved and the need to respect individuals’ IP rights to copyrighted characters, commercial services like DALL·E have begun deploying interventions that prevent generating copyrighted characters. For example, DALL·E rejects user requests that directly name copyrighted characters (e.g., “Mario”). It also rewrites user’s prompts that target specific characters into more generic descriptions (OpenAI, 2024). However, such interventions have never been systematically assessed. Our work fills this gap by making the following key contributions.
@@
First, we build CopyCat , an evaluation suite with a diverse set of popular copyrighted characters and a novel evaluation pipeline (§2), to better ground our investigation and evaluation. Our evaluation captures the inherent trade-off between elimination of similar output and consistency with user intent. We then evaluate five image generation models: Playground v2.5 (Li et al., 2024a), Stable Diffusion XL (Podell et al., 2024), PixArt-$\alpha$ (Chen et al., 2024), DeepFloyd IF (StabilityAI, 2023), and DALL·E 3 (Betker et al., 2023), as well as one video generation model, VideoFusion (Luo et al., 2023).
@@
Figure 1: Examples of copyrighted characters generated by the open-source Playground v2.5 model (Li et al., 2024a) and proprietary DALL·E 3 model. The figures show Mario (a) and Batman (b), which can be generated with their names directly included in the prompt (character name anchoring, though DALL·E 3 rejects the generation with its built-in guardrails with messages like, “I can’t generate an image of Mario/Batman due to content policy restrictions”) or without their names using relevant keywords (indirect anchoring, still possible for DALL·E 3 despite its guardrails).
Second, we demonstrate that as little as two generic keywords associated with a character are enough to generate their image, without mentioning their name (Figure 1). We categorize text that triggers copyrighted character generation into different modes, depending on whether the text explicitly contains character name (Character Name Anchoring), or generic keywords and descriptions only (Indirect Anchoring). We use the term ‘anchor’ to refer to these different sources of generation. Similar to our indirect anchoring concept, some previous work have observed that to obtain certain objects or concepts in generated image, the prompt does not necessarily need to include its name (Zhang et al., 2024; Chin et al., 2024; Lee et al., 2024). Indirect anchoring is particularly crucial for both model deployers and model users: even a non-malicious user could inadvertently generate copyrighted characters when using seemingly innocuous prompts, leading to potential legal liability for the model deployer as well as any unsuspecting user attempting to monetize the generated image.
We systematically evaluate these two modes of generation using CopyCat. We propose a two-stage method to identify indirect anchors: We first use a language model to generate candidate keywords and descriptions that might evoke the likeliness of a copyrighted character. Then we rerank based on embedding space distance to, and co-occurrence with, character names in popular training corpora (§3). Our findings suggest a mismatch between the level of generality in the prompt versus the specificity of the output: For example, $40\%$ of our target characters can be recreated with only 5 automatically selected keywords using our method (§5.1), and some with as few as 2 keywords. These indirect anchors also extend to video generation models and product systems with built-in safeguards, like DALL·E.
@@
Third, we find that existing mitigations are not fully effective and suggest new strategies (§5.2). In this paper we focus on runtime approaches only, assuming that models cannot be modified to remove copyrighted characters. We explore practical solutions that model deployers can incorporate into a production system. We find that prompt rewriting—an important technical piece for copyright protection adopted by model deployers—is far from perfect (§5.2). Applying this intervention alone can only reduce the number of generated characters by half compared to no intervention—and many of the failed rewritten prompts contain indirect anchoring keywords we identified. Instead, combining this strategy with negative prompting (i.e., steering models away from concepts like “red hat”, a defining feature of Mario, during inference) significantly boost mitigation effectiveness. Applying this combination reduces the generation of over 80% of copyrighted characters compared to no mitigation. We also demonstrate that a combination of mitigation strategies can strike a balance between effectively eliminating similar outputs and adhering to user intent.
We summarize the key takeaways for users and model deployers as follows:
@@
  • We call for more awareness of indirect anchoring, where models can generate copyrighted characters without explicitly mentioning the character’s name. For deployers, this may bypass safeguards that rely on direct name detection. For users, such prompts can result in the generation of characters substantially similar to copyrighted ones, leading to potential liability even if they did not intend to generate them.
@@
  • For model deployers who adopt mitigation strategies and intend to prevent the generation of copyrighted characters, we recommend investing in techniques beyond prompt rewriting. Our work suggests a relatively simple set of strategies, such as combining prompt rewriting and negative prompts, to make progress on this issue.
2 CopyCat: An Evaluation Suite for Copyrighted Characters Detection
To systematically study the anchors leading to the generation of copyrighted characters and how to prevent it, we curate a list of copyrighted characters (§2.1) and develop a novel evaluation pipeline (§2.2) that we call CopyCat (Copyrighted CharAcTers).
2.1 A Curation of Copyrighted Characters
We first curate a character list comprising a diverse selection of copyrighted characters to concretely study the effects of character name and indirect anchoring. We source copyrighted characters from popular studios and franchises, as they are more likely to have been present in the training process of image and video generation models. These characters represent a diverse distribution from superhero movies (e.g., Batman, Iron Man, Hulk), animations (e.g., Lightning McQueen, Monkey D. Luffy, Elsa), and video games (e.g., Mario, Pikachu, Link). In addition to U.S. studios like Disney and DreamWorks, we also include international ones like Nintendo and Shogakukan. In total, our collection includes 50 diverse popular copyrighted characters from 18 different studios and subsidiaries. The full list of characters in CopyCat can be found in Appendix B. Throughout this paper, we refer to this curated list of characters as $\mathcal{D}$.
2.2 Evaluation Metrics
In the context of copyright and generative models, especially when any intervention strategy is applied, there is an inherent trade-off between two key factors: 1) dissimilarity from copyrighted entities (e.g., avoiding the generation of specific copyrighted characters like Mario) and 2) consistency with user intent (e.g., if the user requests a plumber, still generating a plumber). Both factors warrant quantification, and an effective intervention method should strike a balance between these two aspects. Specifically, for a copyrighted character $C$ and a corresponding generated image $\mathcal{I}=f_{p,m}(C)$, where $f(\cdot)$ is the generation model, $p$ is the given prompt, and $m$ is mitigation (if any), we calculate the following two metrics:
Detected similarity to copyrighted characters. In general, the more similar a generated image is to existing copyrighted characters, the higher the likelihood of potential legal issues like copyright infringement. In our evaluation, we use GPT-4V (Achiam et al., 2023) to detect whether $C$ is present in $\mathcal{I}$ (see Appendix C). [^2] We use GPT-4V detection for research evaluation purposes, but we are not suggesting that this is a reliable detector for legal judgement of copyright infringement, as such judgments typically require case-by-case analysis. The detector outputs $d(C,\mathcal{I})\in\{0,1\}$, indicating the presence (1) or absence (0) of character $C$ in image $\mathcal{I}$. We then define the metric $\mathsf{DETECT}$ for a model $f$, prompting method $p$, mitigation $m$ as:
@@ 2.2 Evaluation Metrics
which sums the binary detection scores across the character list $\mathcal{D}$. A lower $\mathsf{DETECT}$ score indicates that fewer copyrighted characters were generated. We include human evaluation of copyrighted character detection in Appendix D.
Consistency with user intent. On the other hand, if a model always outputs a random image or rejects the user’s request, it can achieve near-perfect elimination of similar output to copyrighted characters but would fail to fulfill the user’s intent. Therefore, we quantify the consistency between the generation and the user’s intent as a potential metric for user satisfaction: we test whether the key characteristics in user prompt can be found in the generated image. Since we assume that users are prompting the models to obtain certain characters, we can generate the ground-truth key characteristics using the list of target characters. For each copyrighted character $C$ in CopyCat’s curated list of characters, we ask GPT-4 to automatically identify its main general characteristics $s(C)$, which we manually verify and adjust if necessary (e.g., “cartoon mouse” for Mickey Mouse). We then use VQAScore (Lin et al., 2024) to measure the consistency between image $\mathcal{I}$ and characteristics $s(C)$, defined as $c(s(C),\mathcal{I})=\mathbb{P}(\texttt{``Yes''}|\mathcal{I},``\texttt{Does this figure show}~{}s(C)\texttt{? Please answer yes or no.}")$. " ). For example, we calculate (‘‘$\mathbb{P}(\texttt{``Yes''}|\mathcal{I},``\texttt{Does this figure show a cartoon mouse}\texttt{? Please answer yes or no.}")$. " ) when the character is Mickey Mouse. The consistency metric $\mathsf{CONS}$ for a model $f$, input prompt $p$, and intervention $m$ is the average consistency score across the curated list of characters in CopyCat’s character list:
@@
A higher $\mathsf{CONS}$ indicates better consistency with user intent (see Section C.4 for examples). The underlying assumption is that as long as the main general characteristic (e.g., a “cartoon mouse”) is present in the generation, the user may still be more satisfied with the result despite some alterations. We note that this only captures some aspects of consistency with user requests and that future work may improve this metric.
We omit $(f,p,m)$ for $\mathsf{DETECT}$ and $\mathsf{CONS}$ if they are clear from the context. When studying the effect of Character Name and Indirect Anchoring, our evaluation only relies on $\mathsf{DETECT}$, since this task’s focus is whether specific characters can be generated. The trade-off becomes more relevant when involving intervention, hence both $\mathsf{DETECT}$ and $\mathsf{CONS}$ are used for evaluating these interventions: An effective intervention strategy $m$ should aim to minimize $\mathsf{DETECT}$ while maximizing $\mathsf{CONS}$. CopyCat provides a useful framework for understanding different modes of copyrighted character generation and the effectiveness of mitigation strategies, which we will discuss in the following sections.
@@ 3 Identifying Indirect Anchors
Not surprisingly, prompting with “Mario” would likely generate this Nintendo character. We refer to this type of generation as Character Name Anchoring. However, if users ask for a generic “video game plumber” they will also receive the iconic character’s likeness from most models (Figure 1(a)). We refer to this mode of generation, using keywords or descriptions without the character’s name, as Indirect Anchoring. To reduce copyright violation risk coming from the Indirect Anchoring mode in particular, we would like to first generate a set of indirect anchors for a certain character. To systematically identify such anchoring keywords and understand how they become effective triggers, we further use a two-stage approach involving generation and reranking. We generate a set of candidate descriptions and keywords related to the given character and then rerank the candidates to select the most likely anchors that can effectively trigger the desired character generation.
Generation. First, we use GPT-4 to generate a set of candidate descriptions and keywords pertaining to the visual appearance of the characters in CopyCat, using the prompting template in C. [^3] The textual descriptions are around 60 words in length. This length limit provides maximal descriptive information while keeping under the 77 token limit for stable diffusion models (Urbanek et al., 2023).
@@
  • LM-Ranked: For keywords, we also obtain an inherently LM-ranked list as a baseline for comparison. This is achieved by obtaining the top $k$ keywords associated with certain characters using greedy decoding, based on the prompt template provided in Appendix C. Note that the LM may generate words not present in the candidate list, but we maintain $k$ as the same for a fair comparison between LM-Ranked, EmbeddingSim, and Co-Occurrence.
While we focus on keywords re-ranking in later parts of this paper as they provide valuable information for the design of mitigation strategy, we also include relevant analysis on descriptions in §F.3. As a concrete example can be seen in Figure 6 in Section C.5.

### v2

@@ Abstract
footnotetext: Equal contribution. Emails for leading authors: luxihe@cs.princeton.edu, yangsibo@princeton.edu, and swj0419@uw.edu. Our code is available at https://github.com/princeton-nlp/CopyCat.
Recent studies show that image and video generation models can be prompted to reproduce copyrighted content from their training data, raising serious legal concerns about copyright infringement. Copyrighted characters (e.g., Mario, Batman) present a significant challenge: at least one lawsuit has already awarded damages based on the generation of such characters. Consequently, commercial services like DALL·E have started deploying interventions. However, little research has systematically examined these problems: (1) Can users easily prompt models to generate copyrighted characters, even if it is unintentional?; (2) How effective are the existing mitigation strategies? To address these questions, we introduce a novel evaluation framework with metrics that assess both the generated image’s similarity to copyrighted characters and its consistency with user intent, grounded in a set of popular copyrighted characters from diverse studios and regions. We show that state-of-the-art image and video generation models can still generate characters even if characters’ names are not explicitly mentioned, sometimes with only two generic keywords (e.g., prompting with “videogame, plumber” consistently generates Nintendo’s Mario character). We also introduce semi-automatic techniques to identify such keywords or descriptions that trigger character generation. Using this framework, we evaluate mitigation strategies, including prompt rewriting and new approaches we propose. Our findings reveal that common methods, such as DALL·E’s prompt rewriting, are insufficient alone and require supplementary strategies like negative prompting. Our work provides empirical grounding for discussions on copyright mitigation strategies and offers actionable insights for model deployers implementing these safeguards.
@@ 1 Introduction
State-of-the-art image and video generation models demonstrate a remarkable ability for generating high-quality visual content based on free-form user inputs (Rombach et al., 2022; Betker et al., 2023; Chen et al., 2024; Li et al., 2024a; Blattmann et al., 2023; Esser et al., 2024). However, recent research has shown that generative models, including those for image and video, are susceptible to memorizing and generating entire datapoints or concepts from their training data (Somepalli et al., 2023; Carlini et al., 2023; car, 2023). Since some training data originates from copyrighted materials (car, 2023; Kumari et al., 2023), regurgitation of such content may lead to legal intellectual property liability for users and model deployers who further make use of the generated content. In particular, this liability may stem not only from verbatim generation of training data points, but generation of some copyrightable repeating motifs highly similar to those from the training data.
As a result, copyright concerns in image generative models has been extensively discussed in both academic research (car, 2023; Ma et al., 2024; Kim et al., 2024) and litigation (Vincent, 2023; Andersen et al. v. Stability AI et al., N.D. Cal. 2023). Among the diverse copyrighted content that models may generate, copyrighted characters pose a unique legal challenge (Sag, 2023; Henderson et al., 2023; Lee et al., 2024), since characters are a type of “abstractly” protected content (Sag, 2023) Copyrighted characters also pose a unique technical challenge, as they are like general concepts that can appear in many poses, sizes, and variations, making it challenging to apply many previous methods designed for (near) verbatim memorization. As a result, commercial platforms like DALL·E have started implementing tailored measures, such as prompt rewriting, with the goal of preventing the generation of copyrighted characters.
However, two key questions regarding the infringement of copyrighted characters in generative models have not been systematically studied: ( Q1Q1) How easy is it to prompt models to generate copyrighted characters, especially when the copyrighted character is prompted indirectly or generated inadvertently? ( Q2Q2) How effective are common mitigation approaches at reducing copyrighted characters in image generation models?
@@
To address these questions, we present novel evaluation metrics that measure both the generated image’s resemblance to copyrighted characters and its consistency with user input. We use a diverse set of 50 popular characters from various studios and regions as the basis for our analysis. We use this framework to evaluate five image generation models – Playground v2.5 (Li et al., 2024a), Stable Diffusion XL (Podell et al., 2024), PixArt-$\alpha$ (Chen et al., 2024), DeepFloyd IF (StabilityAI, 2023), and DALL·E 3 (Betker et al., 2023) – along with one video generation model, VideoFusion (Luo et al., 2023), using a diverse set of popular copyrighted characters (§ 2).
@@
Figure 1: Examples of copyrighted characters generated by the open-source Playground v2.5 model (Li et al., 2024a) and proprietary DALL·E 3 model. For Mario (a) and Batman (b), both models can generate these characters through indirect anchoring, using relevant descriptive keywords instead of character names. DALL·E 3 blocks explicit name prompts requests (character name anchoring) with content policy messages.
We study Q1Q1 using our evaluation framework, showing that as few as two generic keywords associated with a character are enough to generate their image, without mentioning their name (Fig. 1). We refer to this type of generation as Indirect Anchoring, and introduce a two-stage method to identify effective indirect anchors for both image and video models (§ 3). Our findings suggest a mismatch between the level of generality in the prompt versus the specificity of the output.
@@
In investigating Q2Q2, we find that existing mitigations are not fully effective and suggest new strategies (§ 5.2).***This paper focuses on runtime approaches only, assuming that models cannot be modified to remove copyrighted characters. We explore practical solutions and simulate what model deployers can incorporate into a production system. We find that prompt rewriting is far from perfect, though combining with negative prompting (e.g., steering models away from concepts like “red hat”, a defining feature of Mario) during inference can help strike a balance between effectively eliminating similar outputs and adhering to user intent.
Our findings suggest that generating copyrighted characters, whether intentionally or unintentionally, is relatively easy, and existing mitigation strategies are largely ineffective at preventing this. We thereby summarize the key takeaways for users and model deployers below and hope this work encourages further research in this area.
@@
  • We call for more awareness of indirect anchoring, where models can generate copyrighted characters without explicitly mentioning the character’s name. Deployers should be cautious, as this could bypass safeguards that depend on direct name detection. Users should also be aware that even using generic keywords to unintentionally generate these characters may lead to legal liability.
@@
  • For model deployers who adopt mitigation strategies and intend to prevent the generation of copyrighted characters, relying solely on prompt rewriting is insufficient. We recommend combining prompt rewriting with negative prompts as a more effective approach.
2 An Evaluation Framework for Copyrighted Character Generation
In this section, we describe an evaluation framework to quantify copyrighted character generation and efficacy of a mitigation strategy by combining a metric for copyright protection and consistency with user intent, respectively. Even though our main focus is copyright compliance (e.g., avoiding the generation of specific copyrighted characters like Mario), consistency with user intent is equally important (e.g., if the user requests a plumber, still generating a plumber). A balance between the two factors is usually implicitly considered in real-world intervention strategies such as prompt rewriting (see § 4). To explicitly characterize such inherent trade-off, we define the following metrics.
Copyright protection. As legal judgments of copyright infringement are usually multifaceted and made on a case-by-case basis (see § 1 and § A), it is legally infeasible to have a universal quantitative definition of infringement. Nonetheless, many jurisdictions examine whether two characters are substantially similar (Sag, 2023). Therefore, we study a highly relevant subquestion that can be better quantified: “Is the generated character so similar as to be recognized as an existing copyrighted character?” A “YES” is closely associated with a higher probability of a generation being considered infringing in many jurisdictions. Specifically, for a copyrighted character $C$ and a corresponding generated image $\mathcal{I}=f_{p,m}(C)$, where $f(\cdot)$ is the generation model, $p$ is the given prompt, and $m$ is mitigation (if any), the detector outputs $d(C,\mathcal{I})\in\{0,1\}$, indicating the presence (1) or absence (0) of character $C$ in image $\mathcal{I}$. We then calculate the metric $\mathsf{DETECT}$ by
@@ 2 An Evaluation Framework for Copyrighted Character Generation
which sums the binary detection scores across a character list $\mathcal{D}$ we study. A lower $\mathsf{DETECT}$ score indicates that fewer copyrighted characters were generated. Ideally, a panel of humans would annotate all outputs, but this does not scale. We use GPT-4V as an automatic evaluator for its high accuracy (82.5%) and correlation with humans. We include detector choice ablation and human evaluation of copyrighted character detection in § C.3.
Consistency with user intent. On the other hand, if a model always outputs a random image or rejects requests, it minimizes similarity to copyrighted characters but fails to meet user intent. We quantify consistency between generation and intent as a proxy for user satisfaction, testing whether key prompt characteristics appear in the image. The assumption is that as long as general traits (e.g., a “cartoon mouse”) are present, users may still be satisfied despite alterations. This evaluation reflects how model creators balance legal risks with user experience: tools like DALL·E still respond to user requests but attempt (sometimes unsuccessfully) to steer away from exact recreations of copyrighted characters. Assuming users seek specific characters, we can generate ground-truth characteristics from the target character lists. For each copyrighted character $C$, we ask GPT-4 to automatically identify its main general characteristics $s(C)$, which we manually verify and adjust if necessary (e.g., “cartoon mouse” for Mickey Mouse). We then use VQAScore (Lin et al., 2024) to measure the consistency between image $\mathcal{I}$ and characteristics $s(C)$, defined as $c(s(C),\mathcal{I})=\mathbb{P}(\texttt{``Yes''}|\mathcal{I},``\texttt{Does this figure show}~{}s(C)\texttt{? Please answer yes or no.}")$. " ). For example, we calculate (‘‘$\mathbb{P}(\texttt{``Yes''}|\mathcal{I},``\texttt{Does this figure show a cartoon mouse}\texttt{? Please answer yes or no.}")$. " ) when the character is Mickey Mouse. The consistency metric $\mathsf{CONS}$ for a model $f$, input prompt $p$, and intervention $m$ is the average consistency score across the list of characters $\mathcal{D}$ we study:
@@
A higher $\mathsf{CONS}$ indicates better consistency with user intent. We show that the $\mathsf{CONS}$ evaluator has high agreement with humans in § C.4. We note that current definitions and metrics only capture some aspects of consistency with user requests, and future work can further improve upon this definition.
When evaluating what kind of prompt could trigger the generation of a copyrighted character in an unprotected system, we focus on $\mathsf{DETECT}$. When evaluating a protected system and intervention strategy, the trade-off becomes relevant. In this case, we consider both metrics. An effective intervention strategy $m$ should aim to minimize $\mathsf{DETECT}$ while maximizing $\mathsf{CONS}$. We omit $(f,p,m)$ for $\mathsf{DETECT}$ and $\mathsf{CONS}$ in the paper if they are clear from the context.
We apply this evaluation framework to a set of 50 diverse copyrighted characters, denoted as $\mathcal{D}$ in the paper, to obtain quantitative evaluation on character name, indirect anchoring, and mitigation strategies. They are sourced from popular studios and franchises, both U.S. and international. We cover characters from superhero movies (e.g., Batman, Iron Man, Hulk), animations (e.g., Lightning McQueen, Monkey D. Luffy, Elsa), and video games (e.g., Mario, Pikachu, Link), among others. Full list and details of the curated characters are in § B. We provide additional experiments in § C.3 to show that the evaluator can generalize to a larger set of characters.
In the following sections, we discuss how we apply the evaluation framework to better understand different modes of copyrighted character generation and the effectiveness of mitigation strategies.
@@ 3 Identifying Indirect Anchors
Not surprisingly, prompting with “Mario” would likely generate this Nintendo character. We refer to this type of generation as Character Name Anchoring. However, if users ask for a generic “video game plumber” they will also receive the iconic character’s likeness from most models (Figure 1(a)). We refer to this mode of generation, using keywords or descriptions without the character’s name, as Indirect Anchoring. We propose a systematic method to identify such indirect anchors, revealing their widespread role in generating copyrighted characters (§ 5.1) and informing strategies using indirect anchors to mitigate copyright risks (§ 5.2). Our two-stage approach first generates candidate descriptions and keywords for a character, then reranks them to identify the most effective triggers.
Generation. First, we use GPT-4 to generate a set of candidate descriptions and keywords pertaining to the visual appearance of the characters we apply our evaluation framework on, using the prompting template in C.5.The textual descriptions are around 60 words in length. This length limit provides maximal descriptive information while keeping under the 77 token limit for stable diffusion models (Urbanek et al., 2023).
@@
  • LM-Ranked: For keywords, we also obtain an inherently LM-ranked list as a baseline for comparison. This is achieved by obtaining the top $k$ keywords associated with certain characters using greedy decoding, based on the prompt template provided in § C. Note that the LM may generate words not present in the candidate list, but we maintain $k$ as the same for a fair comparison between LM-Ranked, EmbeddingSim, and Co-Occurrence.
While we focus on keywords re-ranking in later parts of this paper as they provide valuable information for the design of mitigation strategy, we also include relevant analysis on descriptions in § E.3. We provide examples of such keywords and description in Fig. 7 in § C.5.